# Object Re-identification via Spatial-temporal Fusion Networks and Causal Identity Matching

### v1

@@ V Datasets
  • Vehicle-3I is our new dataset for vehicle ReID, as summarized in Table I and Fig. 6. $11$ synchronized cameras captured vehicles at three different intersections. The dataset contains over 40,000 images with 2,038 IDs. The three intersection scenarios reflect the complexity and challenges of real-world traffic environments. Unlike public datasets [13, 14] providing only cropped image patches, it provides full-frame videos. Furthermore, compared to VeRi776 [14], most vehicles in Vehicle-3I are achromatic ($94.21\%$), which poses a significant challenge for ReID (Fig. 7). Achromatic colors without hue, such as white, gray, and black, are difficult to discriminate under varying illumination conditions.
  • Market-1501 [13] contains over 32,000 images of 1,501 different people IDs captured by six non-overlapping synchronized cameras. Each image contains object IDs, timestamps (frame No.), and camera IDs. On average, each person is represented by 3.6 images per viewpoint.
@@ VI Experimental Results
VI-A Settings
To estimate the camera network topology $\mathbb{G}=(\mathbb{V},\mathbb{E})$, we used training datasets (576 IDs in VeRi776, 1,630 IDs in Vehicle-3I and 751 IDs in Market-1501). Among them, 90% of the training data were used for appearance-based ReID model training and the remaining 10% were used for FusionNet training. Note that the object identities (IDs) were completely separated for each training task.
In VeRi776 with 20 cameras ($N_{cam}=20$), the estimated camera network topology contains 400 object transition time distributions $(p_{ij})$. Among them, 380 distributions are between different camera pairs (i.e., $p_{ij}$, where $c_{i}\neq c_{j}$), and each distribution has 300 bins, each covering 100 frame ranges. All distributions were estimated based on the proposed adaptive Parzen window. By performing the same processes, we also estimated the camera network topology of Vehicle-3I and Market-1501.
For the appearance-based ReID model in our framework, we trained FastReID [5] and SOLIDER [19] to extract the appearance similarity. The hyper parameters of FastReID in both datasets are as follows: epoch – 60, batch size – 64. For the vehicle ReID task, a simple ResNet-50 [38] was utilized as the backbone network structure of FastReID. Meanwhile, for the Person ReID task, ResNet-101 with MGN [39] was utilized as the backbone network structure of FastReID.
The proposed FusionNet has a single hidden layer, and we designed the number of nodes in the hidden layer to be around 65% of the size of the input vector by rounding $(2(2W+2)/3+1)$. Training parameters of FusionNet are as follows: epoch – 40, batch size – 128, learning rate – 0.001, optimizer – Adam. VeRi776 [14], Vehicle-3I and Market-1501 [13] have 200, 408, and 750 different IDs for the performance validation of methods and evaluation. We evaluated the rank-$1$ and rank-$5$ accuracy and mean average precision (mAP) based on the traditional 1-to-all single-shot comparison as described in Fig. 4.
@@ VI-A Settings
TABLE II: ReID performances according to $W$ values in FusionNet. The method FastReID [5] does not exploit spatial-temporal information $S_{T}$. FastReID+Wang’s estimates the final similarity by $S_{F}=S_{A}(\textbf{I}^{m}_{i},\textbf{I}^{n}_{j})\cdot p_{ij}\left(\tau^{mn}_{ij}\right)$, as in [10] and employed FastReID [5] for its appearance model. The best performances are marked in bold.
@@ VI-B Effects of FusionNet
In this experiment, we tested on VeRi776 [14], Market-1501 [13], and Vehicle-3I datasets to verify the effect of W value on FusionNet, and the results are shown in Table II. Additionally, we compared other methods such as FastReID [5] using only appearance similarities $(S_{A})$ between images. FastReID + Wang’s estimated the appearance similarity $(S_{A})$ based on FastReID [5] and simply combined the spatial-temporal similarity by $S_{F}=S_{A}(\textbf{I}^{m}_{i},\textbf{I}^{n}_{j})\cdot p_{ij}\left(\tau^{mn}_{ij}\right)$, as in [10].
@@
In VeRi776 [14], the FastReID [5] model, which uses only appearance information, achieved a rank-1 accuracy of 96.96% and an mAP of 81.91On the other hand, combining FastReID with Wang’s approach, which merges additional spatial-temporal similarity, improved the mAP of the appearance-based ReID [5] by 3.56%. However, the simple combination of these two different similarities using $S_{A}(\textbf{I}^{m}_{i},\textbf{I}^{n}_{j})\cdot p_{ij}\left(\tau^{mn}_{ij}\right)$ is not yet optimized. The rank-1 and rank-5 performances decreased after implementing Wang’s approach in FastReID [5]. Using the proposed FusionNet, we achieved a significant improvement in ReID performance. For fair comparisons, we used FastReID [5] as the appearance model for our framework and trained FusionNet with 526 IDs for the appearance model and 50 IDs for FusionNet training. With $W$ set to 10, FusionNet achieved the best performance in rank-1 accuracy at 99.70%, rank-5 accuracy at 99.82%, and mAP at 91.71%. Except at $W=0$, FusionNet outperformed other methods in all evaluation metrics (rank-1, -5, and mAP).
@@
In Vehicle-3I, FastReID [5] using only appearance information achieved a rank-1 accuracy of 63.73% and mAP of 64.24%. The rank-1, rank-5, and mAP performances were increased after adopting wang’s approach in FastReID [5]. This means that using spatial-temporal information in Vehicle-3I is also effective. Except $W=0$, our proposed method significantly outperformed FastReID [5] and FastReID [5]+Wang’s [10]. In $W=6$, we achieved the best performance with 89.22% rank-1 accuracy, 93.50% rank-5 accuracy, and 83.62% mAP.
@@
In Market-1501 [13], which is a person ReID dataset, FastReID [5] achieved a rank-1 accuracy of 96.35% and mAP of 90.77% using only appearance information. However, the rank-1 and mAP performances decreased after adapting Wang’s approach into FastReID [5]. Nevertheless, the proposed framework significantly enhanced the ReID performance. FusionNet achieved the best performance with a rank-1 accuracy of 99.11%, rank-5 accuracy of 99.58%, and mAP of 93.8% when $W$ was set to 10. This result supports that the proposed FusionNet can optimally combine different types of information, such as appearance similarity and spatial-temporal information. In addition, even if the training image for the appearance model was less utilized, it achieved superior performance compared to other methods.
@@
Figure 8: A visualization of the trained weight vector ($\mathbf{w}_{1}$) between input and hidden layer ($\mathbf{h}_{1}$)
@@
Fig. 8 illustrates the trained weight vector ($\mathbf{w}_{1}$) between input and hidden layer ($\mathbf{h}_{1}$). The index of the nodes in $\mathbf{h}_{1}$ is denoted by the row numbers (0–14), and the index of the input vector $\mathbf{x}^{mn}_{ij}$ is denoted by the column numbers (0–21). The first column (0 index) shows the weights for appearance similarity ($S_{A}$), which are notably more significant than other weights, indicating that the FusionNet has been effective in training the importance of appearance similarity ($S_{A}$). The weights in columns 1 to 21 are for spatial-temporal distribution ($S_{T}$). Interestingly, the weights from the 10th to 12th columns showed large magnitudes. This result suggests spatial-temporal information around the time difference ($\tau^{mn}_{ij}$) between two images plays a key role in ReID. Additionally, FusionNet has a lightweight structure but effective.
@@
We also tested the effect of hidden layer depth in FusionNet. We observed that the one hidden layer achieved the best mAP performance (91.71%). The performance decreased significantly to a 59.47% mAP score with a simple perceptron without a hidden layer. This means that a simple perceptron structure is not sufficient for combining appearance similarity and spatial-temporal information. Interestingly, increasing the depth of hidden layers did not lead to an improvement in mAP score; relatively, the mAP scores tended to decrease, E.g., 2-layer: 91.22%, 3-layer: 91.28%, 4-layer: 91.22%. It suggests that a large depth of hidden layers is not necessary, and one hidden layer is sufficient for learning the relationships between appearance similarity and spatial-temporal similarity.
@@
Figure 9: Performance according to $\alpha$ and $\beta$ factors of the adaptive Parzen window.
@@ VI-C Effects of adaptive Parzen window
We evaluated various factors for the proposed adaptive Parzen window method, such as the scale factor $\alpha$ and smoothness factor $\beta$ in Eq. 4. We set $W=10$ for FusionNet on VeRi776 [14] and Market-1501 [13] and set $W=6$ on Vehicle-3I for FusionNet. Fig. 9 illustrates the mAP according to the $\alpha$ and $\beta$ factors. When the scale factor $\alpha=6$, and the smoothness factor $\beta=25$, the proposed framework achieved the best mAP performance on VeRi776 [14]. In the same way, the proposed framework achieved the best mAP performance on all datasets with $\alpha=6,\beta=20$ on Vehicle-3I and $\alpha=4,\beta=120$ on Market-1501 [13]. Note that ReID performances do not significantly change with the factor values, as the difference between the maximum mAP and minimum mAP is 0.09%, 1.77%, and 0.47% in VeRi776 [14], Vehicle-3I and Market-1501 [13].
We further compared the ReID performances for the fixed and proposed adaptive $\sigma$ values. The fixed $\sigma=1,5,10,100$, and the adaptive $\sigma$ as in Eq. 2 were tested. As summarized in Table III, the fixed $\sigma=100$ performed the worst because a too-large $\sigma$ leads to smoothed distributions close to a uniform distribution and reduces the influence of spatial-temporal information. On the other hand, $\sigma=1$, $\sigma=5$ and $\sigma=10$ performed relatively well. However, compared to using fixed $\sigma$ values, the proposed adaptive Parzen window with the adaptive $\sigma$ performed the best in the evaluation metrics (Rank-1, mAP) for all tested datasets. The proposed adaptive Parzen window achieved with 99.70%, 89.22%, 99.11% rank-1 accuracy, and 91.71%, 83.62%, 93.80% mAP on three different datasets, respectively. As explained in Section III-A and Fig. 3, a fixed value of $\sigma$ has difficulty handling various types of initial histograms $h_{ij}$. This result implies that setting different $\sigma_{ij}$ values by considering the various connection strengths of camera pairs is effective and improves ReID performance.
@@
TABLE III: ReID performance according to the values of $\sigma$ for the Parzen window method. The best performances are marked in bold.
@@ VI-D Comparison with state-of-the-art methods
In this section, we compare the proposed method with state-of-the-art re-identification methods using the VeRi776 [14], Vehicle-3I vehicle ReID dataset and Market-1501 [13] person ReID dataset. The methods are primarily categorized into two approaches: 1) only appearance-based, 2) using additional spatial-temporal information (marked by $\dagger$). The methods marked by a ‘*’ performed re-ranking post-processing for the final ReID results.
Table IV summarizes the vehicle ReID results, sorted by rank-1 performance. Our framework achieved outstanding performance in the VeRi776 [14] dataset, performing a rank-1 accuracy of 99.7% and a mAP score of 91.71%. Despite improvements to their baseline methods, the spatial-temporal approaches [14, 8, 30] did not match the performance of other appearance-based methods due to using older appearance models such as SIFT, bag-of-words, and Siamese-CNN for ReID. Furthermore, these methods did not directly estimate the camera network topology or optimize the use of spatial-temporal information.
Many appearance-based methods have significantly improved ReID performance by developing deep learning models. For instance, our baseline appearance model, FastReID [5], achieved a 96.96% rank-1 accuracy and an 81.91% mAP. RPTM [6] notably used the GMS [40] feature matcher and adopted the ResNet-101 [38] structure, demonstrating the superior performance with a 97.3% rank-1 accuracy and an 88.0% mAP. In contrast, despite using a lightweight ResNet-50 structure for the appearance model, our methods outperformed the sophisticated RPTM [6] approach, achieving a 2.4% higher rank-1 accuracy, a 1.42% higher rank-5 accuracy, and a 3.71% higher mAP score. Furthermore, our methods did not perform any re-ranking processes for post-processing.
For Vehicle-3I, several state-of-the-art methods were evaluated. FastReID [5], achieved a 63.73% rank-1 accuracy and 64.24% mAP. On the other hand, CLIP-ReID [21], which achieved the second-best performance on VeRi776, showed the lowest performance on Vehicle-3I because it was trained without viewpoint annotations in this experiment. Adopting Wang’s [10] approach to FastReID [5] achieved 81.37% rank-1 accuracy and 88.60% rank-5 accuracy, proving the effectiveness of spatial-temporal cues. Our method achieved outstanding results, outperforming all other methods with 89.22% rank-1 accuracy and 83.62% mAP. This result showed that using spatial-temporal information with appearance information is more effective in difficult conditions than using appearance information only.
@@
TABLE IV: Performance comparisons on Vehicle ReID datasets. $\dagger$ and * indicate the spatial-temporal approach and re-ranking, respectively. The best and second best performances are marked in bold and underline.
@@
TABLE V: Performance comparisons on the Market- 1501 [13] Person ReID dataset. $\dagger$ indicate the spatial-temporal approach, respectively. The best and second best performances are marked in bold and underline.
@@
Figure 11: A challenging scenario in ReID. Green and red boxes denote true and false matching. A query image did not carry an umbrella, but it was carrying an umbrella in other cameras (gallery). The baseline (FastReID [5]) only relying on appearance information failed to find true matching in the gallery. In contrast, the proposed methods correctly found true matching.
@@
The Person ReID results from the Market-1501 [13] dataset are summarized in Table V. Despite the st-ReID [10] achieving a reasonable rank-1 accuracy of 98.1%, the mAP of 87.6% was lower than that of other state-of-the-art methods. On the other hand, SOLIDER [19] with Swin Transformer [20] achieved good performance with an mAP of 93.9%, but its rank-1 performance of 96.9% was relatively low. For this dataset, we selected two different appearance-based models, FastReID [5] and SOLIDER [19], as the baseline for our framework. As a result, our frameworks achieved the best rank-1 accuracy (99.11%) with FastReID [5], and the best rank-5 accuracy (99.6%) and mAP (95.5%) with SOLIDER [19]. These results imply that the proposed spatial-temporal framework improves both vehicle and person ReID tasks and has the potential to achieve higher performance when it employs a better appearance-based model as its baseline.
@@
Fig. 10 illustrates the qualitative ReID results. We compared the proposed method (Ours) with the baseline method (FastReID [5]), which used only appearance information. The baseline method shows numerous false matches where the appearance closely resembles the query images. The baseline method incorrectly matched ID 470 and 1260 images in Vehicle-3I, which have the same color and model type. However, the proposed method correctly matches most of the same vehicles. In particular, it rarely matched the correct images of the 662-nd vehicle query image in VeRi776 [14] due to many similar black cars. In contrast, the proposed method, thanks to the spatial-temporal information, perfectly matched the correct images at rank-1 to rank-10 under the challenging query and gallery pairs. The baseline model struggled to match correct pairs in the gallery in the person ReID task. For example, ID 342, wearing a stripe yellow shirt, gray shorts, and carrying a black backpack, is easily confused with others wearing similar outfits. ID 1083, wearing a red dress and carrying a shoulder bag, is also very confusing, as another person is wearing a similar red dress with a handbag.
@@
Interestingly, as shown in Fig. 11, the query image is wearing a white shirt and black pants, but it was not wearing an umbrella at that time. Based on the appearance model, it naturally found the images with white shirts and black pants in the gallery. However, the query was wearing an umbrella in other cameras a few minutes later. Under this challenging ReID scenario, the appearance-based model only relies on the appearance of the query and never finds correct matches. In contrast, our method using spatial-temporal information correctly matched the images, even if the appearance of the query has changed significantly. These results support that the proposed ReID (Ours) based on the spatial-temporal fusion network can effectively manage the appearance ambiguity problems and overcome the limitations of the previous ReID methods.

### v2

@@ V Datasets
  • Vehicle-3I is our new dataset for vehicle ReID, as illustrated in Fig. 6. It includes data from $11$ synchronized cameras capturing vehicles at three intersections. The dataset contains over 40,000 images with 2,038 IDs. These three intersection scenarios reflect the complexity and challenges of real-world traffic environments. Unlike public datasets [13, 14] that provide only cropped image patches, the Vehicle-3I dataset offers full-frame videos. Furthermore, compared to the VeRi776 [14], most vehicles in Vehicle-3I are achromatic ($94.21\%$), which poses a significant challenge for ReID (Fig. 7). Achromatic colors without hue, such as white, gray, and black, are difficult to discriminate under varying illumination conditions.
  • Market-1501 [13] dataset contains over 32,000 images of 1,501 different person IDs captured by six non-overlapping synchronized cameras. Each image includes object IDs, timestamps (frame No.), and camera IDs.
@@ VI Experimental Results
TABLE I: ReID performances according to different $W$ values in FusionNet. The FastReID [5] method does not utilize spatial-temporal information $S_{T}$. FastReID+Wang’s method estimates the final similarity by $S_{F}=S_{A}(\textbf{I}^{m}_{i},\textbf{I}^{n}_{j})\cdot p_{ij}\left(\tau^{mn}_{ij}\right)$, as in [10] and employed FastReID [5] for its appearance model. The best performances are marked in bold.
VI-A Settings
To estimate the camera network topology $\mathbb{G}=(\mathbb{V},\mathbb{E})$, we used training datasets containing  (576 IDs from VeRi776, 1,630 IDs from Vehicle-3I, and 751 IDs from Market-1501). Of these datasets, 90% of the training data was used for appearance-based ReID model training, while the remaining 10% was reserved for FusionNet training. It is important to note that the object identities (IDs) were completely separated for each training task.
In VeRi776, with 20 cameras ($N_{cam}=20$), the estimated camera network topology contains 400 object transition time distributions $(p_{ij})$. Of these, 380 distributions are between different camera pairs (i.e., $p_{ij}$, where $c_{i}\neq c_{j}$), with each distribution consisting of 300 bins, each covering a range of 100 frames. All distributions were estimated using the proposed adaptive Parzen window method. The same processes were applied to estimate the camera network topology for Vehicle-3I and Market-1501.
For the appearance-based ReID model in our framework, we trained FastReID [5] and SOLIDER [19] to extract appearance similarity. The hyperparameters for FastReID across both datasets were as follows: epoch —- 60, batch size -— 64. For the vehicle ReID task, a simple ResNet-50 [38] was utilized as the backbone network structure of FastReID. Meanwhile, for the Person ReID task, ResNet-101 with MGN [39] was utilized as the backbone network structure of FastReID.
The proposed FusionNet has a single hidden layer, with the number of nodes in the hidden layer set to approximately 65% of the input vector size, calculated as $(2(2W+2)/3+1)$ and rounded accordingly. The training parameters for FusionNet are as follows: epoch —- 40, batch size —- 128, learning rate —- 0.001, optimizer —- Adam. The datasets VeRi776 [14], Vehicle-3I and Market-1501 [13] contain 200, 408, and 750 different IDs, respectively, for method validation and performance evaluation. We evaluated rank-$1$ and rank-$5$ accuracy, and as mAP, based on the traditional 1-to-all single-shot comparison, as illustrated in Fig. 4.
@@ VI-B Effects of FusionNet
In this experiment, we tested the effect of the $W$ value on FusionNet using the VeRi776 [14], Market-1501 [13], and Vehicle-3I datasets. The results are presented in Table I. Additionally, we compared other methods, such as FastReID [5], which uses only appearance similarities $(S_{A})$ between images. FastReID + Wang’s method estimates the appearance similarity $(S_{A})$ based on FastReID [5] and combines it with spatial-temporal similarity using $S_{F}=S_{A}(\textbf{I}^{m}_{i},\textbf{I}^{n}_{j})\cdot p_{ij}\left(\tau^{mn}_{ij}\right)$, as described in [10].
@@
In the VeRi776 [14], the FastReID [5] model, which uses only appearance information, achieved a rank-1 accuracy of 96.96% and an mAP of 81.91However, combining FastReID with Wang’s approach, which merges additional spatial-temporal similarity, improved the mAP of the appearance-based ReID [5] by 3.56%. Despite this improvement in mAP, the simple combination of these two different similarities using $S_{A}(\textbf{I}^{m}_{i},\textbf{I}^{n}_{j})\cdot p_{ij}\left(\tau^{mn}_{ij}\right)$ is not yet optimized. As a result, the rank-1 and rank-5 performances decreased after implementing Wang’s approach in FastReID [5]. Using the proposed FusionNet, we achieved a significant improvement in ReID performance. For fair comparisons, we used FastReID [5] as the appearance model in our framework and trained FusionNet with 526 IDs for the appearance model and 50 IDs for FusionNet training. With $W$ set to 10, FusionNet achieved the best performance in rank-1 accuracy at 99.70%, rank-5 accuracy at 99.82%, and mAP at 91.71%. Except at $W=0$, FusionNet outperformed other methods in all evaluation metrics (rank-1, -5, and mAP).
@@
In the Vehicle-3I dataset, FastReID [5] using only appearance information, achieved a rank-1 accuracy of 63.73% and an mAP of 64.24%. The rank-1, rank-5, and mAP performances increased after adopting wang’s approach in FastReID [5]. This indicates that using spatial-temporal information in the Vehicle-3I dataset is also effective. Except for $W=0$, our proposed method significantly outperformed FastReID [5] and FastReID [5]+Wang’s [10]. At $W=6$, we achieved the best performance with 89.22% rank-1 accuracy, 93.50% rank-5 accuracy, and 83.62% mAP.
@@
In the Market-1501 [13], person ReID dataset, FastReID [5] achieved a rank-1 accuracy of 96.35% and an mAP of 90.77% using only appearance information. However, the rank-1 accuracy and mAP decreased after adapting Wang’s approach into FastReID [5]. Despite this, the proposed framework significantly enhanced ReID performance. FusionNet achieved the best performance with a rank-1 accuracy of 99.11%, rank-5 accuracy of 99.58%, and mAP of 93.8% when $W$ was set to 10. These results demonstrate that the proposed FusionNet can optimally combine different types of information, such as appearance similarity and spatial-temporal information. Additionally, even with the reduced training images for the appearance model, it achieved superior performance compared to other methods.
@@
Figure 8: A visualization of the trained weight vector ($\mathbf{w}_{1}$) between the input layer and the hidden layer ($\mathbf{h}_{1}$)
@@
Fig. 8 illustrates the trained weight vector ($\mathbf{w}_{1}$) between the input layer and the hidden layer ($\mathbf{h}_{1}$). The index of the nodes in $\mathbf{h}_{1}$ is denoted by the row numbers (0–14), and the index of the input vector $\mathbf{x}^{mn}_{ij}$ is denoted by the column numbers (0–21). The first column (index 0) represents the weights for appearance similarity ($S_{A}$), which are notably more significant than other weights. This indicates that FusionNet has effectively learned the importance of appearance similarity ($S_{A}$). The weights in columns 1 to 21 correspond to the spatial-temporal distribution ($S_{T}$). Interestingly, the weights in columns 10 to 12 exhibit large magnitudes. This result indicates that spatial-temporal information related to the time difference ($\tau^{mn}_{ij}$) between two images plays a key role in ReID. Additionally, FusionNet has a lightweight yet effective structure.
@@
We also tested the effect of hidden layer depth in FusionNet. We observed that a single hidden layer achieved the best mAP performance (91.71%). The performance decreased significantly to a 59.47% mAP score with a simple perceptron without a hidden layer. This indicates that a simple perceptron structure is insufficient for effectively combining appearance similarity and spatial-temporal information. Interestingly, increasing the depth of hidden layers did not lead to an improvement in mAP score; in fact, the mAP scores tended to decrease: for example, 2-layer: 91.22%, 3-layer: 91.28%, and 4-layer: 91.22%. It suggests that many hidden layers are unnecessary, and a single hidden layer is sufficient for learning the relationships between appearance similarity and spatial-temporal similarity.
@@
Figure 9: Performance according to the $\alpha$ and $\beta$ factors of the adaptive Parzen window.
@@ VI-C Effects of adaptive Parzen window
We evaluated various factors for the proposed adaptive Parzen window method, including the scale factor $\alpha$ and smoothness factor $\beta$ in Eq. 4. We set $W=10$ for FusionNet on the VeRi776 [14] and Market-1501 [13] datasets, and $W=6$ for FusionNet on the Vehicle-3I dataset. Fig. 9 illustrates the mAP according to the $\alpha$ and $\beta$ factors. When the scale factor $\alpha=6$ and the smoothness factor $\beta=25$, the proposed framework achieved the best mAP performance on VeRi776 [14]. Similarly, the proposed framework achieved the best mAP performance with $\alpha=6,\beta=20$ on Vehicle-3I and $\alpha=4,\beta=120$ on Market-1501 [13]. Notably, ReID performances do not significantly change with variations in the factor values, as the difference between the maximum and minimum mAP is 0.09%, 1.77%, and 0.47% for =VeRi776 [14], Vehicle-3I and Market-1501 [13].
We compared the ReID performances using fixed and the proposed adaptive $\sigma$ values. The fixed $\sigma=1,5,10,100$, and the adaptive $\sigma$ as in Eq. 2 were tested. As summarized in Table II, the fixed $\sigma=100$ performed the worst because a large $\sigma$ leads to overly smoothed distributions that approximate a uniform distribution, thereby reducing the influence of spatial-temporal information. Conversely, $\sigma=1$, $\sigma=5$, and $\sigma=10$ performed relatively well. However, compared to the fixed $\sigma$ values, the proposed adaptive Parzen window with the adaptive $\sigma$ achieved the best performance across all evaluation metrics (Rank-1, mAP) for all tested datasets. The proposed adaptive Parzen window achieved rank-1 accuracy of 99.70%, 89.22%, and 99.11%, and mAP scores of 91.71%, 83.62%, and 93.80% on three different datasets, respectively. As explained in Section III-A and shown in Fig. 3, a fixed $\sigma$ value struggles to handle various types of initial histograms $h_{ij}$. This result implies that setting different $\sigma_{ij}$ values by considering the varying connection strengths of camera pairs is effective and improves ReID performance.
@@
TABLE II: ReID performance according to the values of $\sigma$ for Parzen window. The best performances are marked in bold.
@@ VI-D Comparison with state-of-the-art methods
In this section, we compare the proposed method with state-of-the-art re-identification methods using the VeRi776 [14], Vehicle-3I vehicle ReID dataset and Market-1501 [13] person ReID dataset. The methods are primarily categorized into two approaches: 1) appearance-based only and 2) using additional spatial-temporal information (marked by $\dagger$). The methods marked by a ‘*’ performed re-ranking post-processing for the final ReID results.
Table III summarizes the vehicle ReID results, sorted by rank-1 performance. Our framework achieved outstanding performance on the VeRi776 [14] dataset, with a rank-1 accuracy of 99.7% and an mAP score of 91.71%. Despite improvements to their baseline methods, the spatial-temporal approaches [14, 8, 30] did not match the performance of other appearance-based methods due to their use of older appearance models such as SIFT, bag-of-words, and Siamese-CNN for ReID. Additionally, these methods did not directly estimate the camera network topology or optimize the use of spatial-temporal information.
Many appearance-based methods have significantly improved ReID performance by developing deep learning models. For instance, our baseline appearance model, FastReID [5], achieved a 96.96% rank-1 accuracy and 81.91% mAP. RPTM [6] notably used the GMS [40] feature matcher and adopted the ResNet-101 [38] structure, demonstrating superior performance with a 97.3% rank-1 accuracy and an 88.0% mAP. In contrast, despite using a lightweight ResNet-50 structure for the appearance model, our methods outperformed the sophisticated RPTM [6] approach, achieving a 2.4% higher rank-1 accuracy, a 1.42% higher rank-5 accuracy, and a 3.71% higher mAP score. Furthermore, our methods did not perform any re-ranking processes for post-processing.
For Vehicle-3I, several state-of-the-art methods were evaluated. FastReID [5], achieved a 63.73% rank-1 accuracy and 64.24% mAP. In contrast, CLIP-ReID [21], which achieved the second-best performance on VeRi776, showed the lowest performance on Vehicle-3I because it was trained without viewpoint annotations in this experiment. Adopting Wang’s [10] approach to FastReID [5] improved performance to 81.37% rank-1 accuracy and 88.60% rank-5 accuracy, demonstrating the effectiveness of spatial-temporal cues. Our method achieved outstanding results, outperforming all other methods with 89.22% rank-1 accuracy and 83.62% mAP. This result demonstrates that combining spatial-temporal information with appearance information is more effective than using appearance information alone in challenging conditions.
@@
TABLE III: Performance comparisons on Vehicle ReID datasets. $\dagger$ and * indicate the spatial-temporal approach and re-ranking, respectively. The best and second best performances are marked in bold and underline.
@@
TABLE IV: Performance comparisons on the Market- 1501 [13] Person ReID dataset. $\dagger$ indicate the spatial-temporal approach, respectively. The best and second best performances are marked in bold and underline.
@@
The Person ReID results from the Market-1501 [13] dataset are summarized in Table IV. Despite the st-ReID [10] achieving a rank-1 accuracy of 98.1%, its mAP of 87.6% was lower than that of other state-of-the-art methods. In contrast, SOLIDER [19], which uses the Swin transformer [20], performed well with an mAP of 93.9%, though its rank-1 accuracy was relatively lower at 96.9%.. For this dataset, we selected two different appearance-based models, FastReID [5] and SOLIDER [19], as baselines for our framework. Our frameworks achieved the best rank-1 accuracy (99.11%) with FastReID [5], and the best rank-5 accuracy (99.6%) and mAP (95.5%) with SOLIDER [19]. These results imply that the proposed spatial-temporal framework improves both vehicle and person ReID tasks and has the potential to achieve higher performance when it employs a better appearance-based model as its baseline.
@@
Fig. 10 illustrates the qualitative ReID results. We compared the proposed method (Ours) with the baseline method (FastReID [5]), which used only appearance information. The baseline method shows numerous false matches where the appearance closely resembles that of the query images. For instance, the baseline incorrectly matched ID 470 and 1260 images in the Vehicle-3I, which have the same color and model type. In contrast, the proposed method correctly matched most of the same vehicles. Notably, it rarely matched the correct images of the 662-nd vehicle query image in VeRi776 [14] despite many similar black cars. The proposed method, leveraging spatial-temporal information, perfectly matched the correct images at rank-1 to rank-10 under the challenging query and gallery pairs. Conversely, the baseline model struggled to match correct pairs in the gallery during the person ReID task. For example, ID 342, wearing a stripe yellow shirt, gray shorts, and a black backpack, was easily confused with others in similar outfits. Similarly, ID 1083, wearing a red dress and carrying a shoulder bag, was often mistaken for another person wearing a similar red dress with a handbag.
@@
Interestingly, as shown in Fig. 11, the query image depicts a person wearing a white shirt and black pants, but not carrying an umbrella at that time. Based on the appearance model, it naturally found the images with white shirts and black pants in the gallery. However, the query carried an umbrella in other cameras a few minutes later. In this challenging ReID scenario, the appearance-based model, relying solely on visual cues, failed to find the correct matches. In contrast, our method, which incorporates spatial-temporal information, correctly matched the images despite substantial changes in the query’s appearance. These results demonstrate that the proposed ReID method, based on the spatial-temporal fusion network, effectively manages appearance ambiguity and overcomes the limitations of previous ReID methods.
Figure 11: A challenging scenario in ReID. Green and red boxes denote true and false matches. In this case, a query image did not show the subject carrying an umbrella, whereas the subject carried an umbrella in other images captured by different cameras (gallery). The baseline method (FastReID [5]), which relies solely on appearance information, failed the correct match in the gallery. In contrast, the proposed methods successfully identified the true match.